OpenStudy (anonymous):

Solve the system of equations and enter the solution as an ordered pair. 3x + 4y = 38 5x - 5y = -30

OpenStudy (anonymous):

2, 8

OpenStudy (anonymous):

5x - 5y = -30 x - y = -6 x = y - 6 3(y - 6) + 4y = 38 3y - 18 + 4y = 38 7y - 18 = 38 7y = 56 y = 8 x = 8 - 6 = 2 (2, 8)